OpenStudy (anonymous):

what are the solutions of the equation (2x +5)(x-4) = 0

OpenStudy (amistre64):

ab=0 implies that a=0, or b=0, or both are zero

OpenStudy (amistre64):

so... when does: 2x+5 = 0 and when does: x-4 = 0

OpenStudy (anonymous):

the answer is x= 5/2 or x=4 but im not sure how i got that answer

jimthompson5910 (jim_thompson5910):

zoeydenmark, one of those answers you got is incorrect

OpenStudy (anonymous):

no its right my teacher gave me the answer but im not sure how to get it

jimthompson5910 (jim_thompson5910):

2x+5 = 0 2x = -5 x = -5/2 <--- one solution ----------- x-4 = 0 x = 4 <---- the other solution

jimthompson5910 (jim_thompson5910):

So there must be a typo

OpenStudy (anonymous):

yes i meant to put the negative

jimthompson5910 (jim_thompson5910):

oh ok

OpenStudy (anonymous):

sorry